OpenStudy (anonymous):

In a change jar, Tom has nickels and dimes. The number of nickels is two times the number of dimes and the total amount of money in the jar is $1.20. How many dimes are in the jar?

OpenStudy (wolf1728):

n = 2*d .05n + .10d = $1.20 d = n/2 substituting this into the above equation we get, .05n + .10n/2 = 1.20 .1n + .1n = 2.40 .2n = 2.40 n = 12 d = 6